OpenStudy (anonymous):

determine algebraically whether the function is even, odd, or neither even nor odd f(x)=3x^5-7x^3

OpenStudy (anonymous):

before we do algebra, take a look at the exponents what do you see?

OpenStudy (anonymous):

clear or no?

OpenStudy (cherylim23):

Recall that definition of an odd function means that f(x) = -f(x) and the definition of an even function means that f(x) = f(-x).

OpenStudy (anonymous):

hmm if \(f(x)=-f(x)\) that would mean \(f\) is always zero the definition of "odd" is \(f(-x)=-f(x)\)

OpenStudy (cherylim23):

Typographical error there, thanks for correcting